Document Page: First | Prev | Next | All | Image | This Release | Search

File: 072496_may96_decls12_0001.txt
Page: 0001
Total Pages: 1

Subject: COMMANDERS COMMENTS FOR 29 OCT 90                               

Unit: 24TH ID     

Parent  Organization: XVIII CORPS 

Box ID: BX001425

Folder Seq #:        283

Document Number:         33









        COMMANDERS COMMENTS FOR 27 OCT 90

        IST BDE: NONE

        2ND BDE: TRAINING HAS BEEN REVISED TO REFLECT A TEN DAY DELAY IN
        THE OPENING OF SMALL ARMS RANGES. WORK WILL CONTINUE TO IMPROVE
        RANGES DURING PERIOD. DIAHRREA/VOMITTING CONTINUES TO AFFECT
        FORCE. 1-64 AR CANCELLED TRAINING TO FOCUS ALL EFFORTS ON FIELD
        SANITATION AND PERSONNEL HYGIENE IN ATTEMPT TO HALT OUTBREAK.

        197 BDE: SEE ATTACHED


        AVN BDE: AVN BDE CONTINUES TO FOCUS ON NIGHT AIRCRAFT AND VEH
        MAINT. 1-24 AVN ORIENTATION IS ON NIGHT TRAINING. 3-24 AVN
        PROVIDE GS TO DIVISION AND BRIEFED CG 09 OH-58D CAPABILITIES.
        AVN BDE IS PLANNING ON HELLFIRE SHOOT AT FAYSAL RANGE ON I NOV
        90. 1-24 AVN WILL CONDUCT RIGHT COMBINED ARMS ATTACK TRAINNG WIH
        IST BDE ON 29 OCT 90.


        DIVARTY: 1-5 ADA S3 MET WITH DIVARTY S3 TO ADDRESS SHORTFALL IN
        AIR DEFENSE FOR 18TH FA BDE. WE ARE WORXING TO PROVIDE COVERAGE
        FOR THEM WITH EXISTING ASSETS WHILE CDR, 1-5 ADA AND 18TH FA BDE
        REQUEST ADDITIONAL ASSETS FROM XVIII CORPS. DIVARTY S3 WILL
        BRIEF GDP TO 4-41 FA OFFICERS TOMORROW AT 1000.


        2-4 CAV: NO CHANGE FROM YESTERDAY


        DISCOM: THE TEXTRON/CO D MI COMPONENT REBUILD TEAM REPAIRED 3
        FUPPS, ALL OF WHICH WENT TO 24 FSB. CO C CL IX PLT RECEIVED 200
        EA M1008 SAND TIRES FOR 24 SIG BN.

        1-5 ADA: SPENT ALL DAY (26 OCT) IN COORDINATION WITH 11TH ADA
        BDE AND 2-44 ADA (IOIST ADA UNIT). MOST PROBLEMS RESOLVED, 11TH
        ADA PLANS NO ADA SPPORT FOR ISTH FA BDE. I WORKED TO GET THEM TO
        RECONSIDER. MOST PROBLEMS ARE THE RESORT OF IMPROPER OPLAN.

        3D ENG: B/3 ENG IS 90% COMPLETE ON CALFEX RANGE TOWER AND 95%
        COMPLETE ON BLEACHERS WAITING ON BOLTS TO SECURE WOOD TO FRAME.
        THE ROAD INTO CALFEK SITE WILL REQUIRE CONTINUOUS MAINT. A/3ENG
        IS IN STAND-DOWN TODAY. C/3ENG AND B/27TH ENO ARE WORKING ON
 COMPLETED
        90% OF ROADS FOR IST BDE BSA.


        124 MI: A,C, AND TM DRAGON CONDUCTING MF INTERCEPT. CI ESTAB
        LIAISON WITH LOCALS. TRANSLATOR ON CALL TO 3 ENG, 197 EDE AND
        724TH MSB. TWO TRANSLATORS SUPPORTING 3D TRANS, 1ST BDE AND 24TH
        MP'S. TEAM TRAINING CONTINUES FOR TRAILBLAZER AND MSQ 103.

        24TH SIG: CONT TO SUPPORT DIV COMS SYSTEMS.

Document Page: First | Prev | Next | All | Image | This Release | Search


Document 1 f:/Week-20/BX001425/DUTY LOG 28 OCT 90 24 ID M D-MAIN/commanders comments for 29 oct 90:07199610202476
Control Fields 17
File Room = may96_declassified
File Cabinet = Week-20
Box ID = BX001425
Unit = 24TH ID
Parent Organization = XVIII CORPS
Folder Title = DUTY LOG 28 OCT 90 24 ID M D-MAIN
Folder Seq # = 283
Subject = COMMANDERS COMMENTS FOR 29 OCT 90
Document Seq # = 76
Document Date =
Scan Date =
Queued for Declassification = 01-JAN-1980
Short Term Referral = 01-JAN-1980
Long Term Referral = 01-JAN-1980
Permanent Referral = 01-JAN-1980
Non-Health Related Document = 01-JAN-1980
Declassified = 19-JUL-1996